# QN-Mixer: A Quasi-Newton MLP-Mixer Model for Sparse-View CT Reconstruction

### v1

@@ 2.1 Background
${\bm{y}}={\bm{A}}{\bm{x}}+\epsilon,$ (1)
where ${\bm{x}}\in\mathbb{R}^{n}$ is the (unknown) object to reconstruct with $n=h\times w$, ${\bm{y}}\in\mathbb{R}^{m}$ is the data (i.e. sinogram), where $m=n_{v}\times n_{d}$, $n_{v}$ and $n_{d}$ denote the number of projection views and detectors, respectively. ${\bm{A}}\in\mathbb{R}^{n\times m}$ is the forward model (i.e. discrete Radon transform [37]), and $\epsilon\in\mathbb{R}^{m}$ is the system noise. The goal of CT image reconstruction is to recover the (unknown) object, ${\bm{x}}$, from the observed data ${\bm{y}}$. As the problem is ill-posed due to the missing data, the linear system in Eq. 1 becomes underdetermined and may have infinite solutions. Hence, reconstructed images suffer from artifacts, blurring, and noise. To address this issue, iterative reconstruction algorithms are utilized to minimize a regularized objective function with a $L^{2}$ norm constraint:
@@ 2.3 Advancements in Deep Unrolling Networks
       ${\bm{s}}_{t}\leftarrow-{\bm{H}}_{t}\nabla_{{\bm{x}}}J({\bm{x}}_{t})$ ${\bm{x}}_{t+1}\leftarrow{\bm{x}}_{t}+{\bm{s}}_{t}$ ${\bm{z}}_{t}\leftarrow\nabla_{{\bm{x}}}J({\bm{x}}_{t+1})-\nabla_{{\bm{x}}}J({\bm{x}}_{t})$ $\rho_{t}\leftarrow 1/({\bm{z}}_{t}^{\mathsf{T}}{\bm{s}}_{t})$ ${\bm{H}}_{t+1}\leftarrow({\bm{I}}-\rho_{t}{\bm{s}}_{t}{\bm{z}}_{t}^{\mathsf{T}}){\bm{H}}_{t}({\bm{I}}-\rho_{t}{\bm{z}}_{t}{\bm{s}}_{t}^{\mathsf{T}})+\rho_{t}{\bm{s}}_{t}{\bm{s}}_{t}^{\mathsf{T}}$
@@ 3.2 Regularization term: Incept-Mixer
Figure 3: Architecture of our regularization block. It is referred to as “Incept-Mixer” and denoted as ${\mathcal{G}}$ in Eq. 4
@@
$\begin{split}\mathrm{MixerLayer}({\bm{e}}_{t})&=\mathrm{Mix}(\mathrm{MLP}_{\texttt{channel}},\\ &\mathrm{Mix}(\left[\mathrm{MLP}_{\texttt{height}},\mathrm{MLP}_{\texttt{width}}\right],{\bm{e}}_{t}),\end{split}$ (6)
where $\mathrm{Mix}(\mathrm{Layer},{\bm{e}}_{t})=\mathrm{Layer}(\mathrm{LN}({\bm{e}}_{t}))+{\bm{e}}_{t}$, with $\mathrm{LN}$ as layer normalization. $\mathrm{MLP}_{\texttt{height}}$, $\mathrm{MLP}_{\texttt{width}}$ are applied to height and width features, respectively, and $\mathrm{MLP}_{\texttt{channel}}$ to rows and shared. Finally, after $N$ such mixer layers, the regularized sample is transformed back to an image through a patch expansion step to obtain ${\mathcal{G}}({\bm{x}}_{t})$. Consequently, the iterative optimization solution is as follows:
@@ 3.4 Proposed algorithm of QN-Mixer
       ${\bm{s}}_{t}\leftarrow-{\bm{H}}_{t}{\bm{r}}_{t}$ ${\bm{x}}_{t+1}\leftarrow{\bm{x}}_{t}+\mathcal{D}({\bm{s}}_{t})$ ${\bm{r}}_{t+1}\leftarrow\mathcal{E}(\nabla_{{\bm{x}}}J({\bm{x}}_{t+1}))$ ${\bm{z}}_{t}\leftarrow{\bm{r}}_{t+1}-{\bm{r}}_{t}$ $\rho_{t}\leftarrow 1/({\bm{z}}_{t}^{\mathsf{T}}{\bm{s}}_{t})$ ${\bm{H}}_{t+1}\leftarrow({\bm{I}}-\rho_{t}{\bm{s}}_{t}{\bm{z}}_{t}^{\mathsf{T}}){\bm{H}}_{t}({\bm{I}}-\rho_{t}{\bm{z}}_{t}{\bm{s}}_{t}^{\mathsf{T}})+\rho_{t}{\bm{s}}_{t}{\bm{s}}_{t}^{\mathsf{T}}$
@@ 4.1 Experimental Setup
Training details.  For each set of $n_{v}$ views, we train our model for $50$ epochs using $4$ Nvidia Tesla V100 (32GB RAM). We employ the AdamW optimizer [29] with a learning rate of $1\times 10^{-4}$, weight decay $1\times 10^{-2}$, and utilize the mean squared error loss with a batch size of $1$. Additionally, we incorporate a learning rate decay factor of $0.1$ after $40$ epochs. Unrolling iterations for QN-Mixer are set to $T=14$. Incept-Mixer uses a patch size of $p=4$, $d=96$ embedding dimension, and $N=2$ mixer layers. The inverse Hessian size is $64^{2}\times 64^{2}$ with $k=2$ downsampling blocks. Following [50], ${\bm{A}}^{\dagger}$ is implemented using the FBP algorithm for the pseudo-inverse of ${\bm{A}}$.
Evaluation metrics.  Consistent with standard evaluation practices [1, 45, 50], we use the structural similarity index measure (SSIM) with parameters set to: level${~{}\text{=}~{}}5$, a Gaussian kernel of size $11$, and a standard deviation of $1.5$ as our primary performance indicator. Additionally, we complement our evaluation with the peak signal-to-noise ratio (PSNR).
@@ 4.2 Comparison with state-of-the-art methods
In medical imaging, where training data predominantly consists of normal patient images, it is imperative to develop methods that generalize to scans with lesions or anomalies. To address this challenge, we conducted an experiment to evaluate model performance on out-of-distribution (OOD) textures. The frozen models were tested with CT images featuring a randomly positioned white circle with no noise added to the sinograms, as illustrated in the third row of Fig. 4. In Tab. 3, QN-Mixer attains state-of-the-art results across all $n_{v}$ views. First-order unrolling networks such as LEARN and RegFormer exhibit significant PSNR degradation of $-3.1$ dB and $-4.22$ dB, respectively, for $n_{v}=32$, while our method demonstrates a milder degradation of $-2.67$ dB.
Visual comparison.  As it can be seen on Fig. 4 and Fig. 5, FBPConvNet and DuDoTrans exhibit noticeable blurry images with severe artifacts when $n_{v}=32$. While Learned PD and LEARN show satisfactory performance, they struggle with intricate details, like in the liver and spine. In contrast, RegFormer produces high-quality images but faces challenges in generalizing to OOD data. QN-Mixer excels in producing high-quality images with fine details, even under challenging conditions such as $n_{v}=32$ views and OOD data.
Method #Iters Epoch time (s) #Params (M) Memory (GB)
FBPConvNet [17] - 68 31.1 1.30
DuDoTrans [45] - 92 15.0 1.38
Learned PD [1] 10 82 0.25 0.81
LEARN [6] 30 780 4.50 1.85
RegFormer [50] 18 700 5.00 10.19
QN-Mixer (ours) 14 594 8.50 7.83
@@
[…]
@@ 4.3 Ablation Study
Inverse Hessian approximation size.  The results in Tab. 5 emphasize the significant impact of the inverse Hessian approximation size on our performance. When too small, a notable degradation is observed (e.g., $8^{2}\times 8^{2}$), while larger sizes result in performance improvements as the approximation approaches the full inverse Hessian. However, exceeding $64^{2}\times 64^{2}$ was unfeasible in our experiments due to increasing memory costs and memory constraints.
Figure 6: Ablation on the number of unrolling iterations. We compare our QN-Mixer against RegFormer. Left: PSNR (dB); Right: SSIM (%)
@@
Regularization term.  We evaluate the impact of the regularization term in our framework. Our Incept-Mixer is compared against various learned alternatives, including the Inception block [42] and MLP-Mixer block [43]. Additionally, we explore the use of the pseudo-inverse ${\bm{A}}^{\dagger}$ instead of the FBP. The results presented in Tab. 6 show that employing the pseudo-inverse results in a less pronounced degradation ($-0.57$ dB and $-0.28$%), enhancing the interpretability of QN-Mixer. Finally, we test our Incept-Mixer in the first-order framework, highlighting the significance of the second-order latent BFGS approximation with a significant improvement ($+2.06$ dB and $+1.86\%$).
@@ 5 Conclusion
In this paper, we investigate the application of deep second-order unrolling networks for tackling imaging inverse problems. To this end, we introduce QN-Mixer, a quasi-Newton inspired algorithm where a latent BFGS method approximates the inverse Hessian, and our Incept-Mixer serves as the non-local learnable regularization term. Extensive experiments confirm the successful sparse-view CT reconstruction by our model, showcasing superior performance with fewer iterations than state-of-the-art methods. In summary, this research offers a fresh perspective that can be applied to any iterative reconstruction algorithm. A limitation of our work is the memory requirements associated with quasi-Newton algorithm. We introduced a memory-efficient alternative by projecting the gradient to a lower dimension, successfully addressing the CT reconstruction problem. However, its applicability to other inverse problems may be limited. In future work, we aim to extend our approach to handle larger Hessian sizes, broadening its application to a range of problems.
@@
1 Inverse Hessian approximation
@@ 1.1 Approximation via optimization
${\bm{H}}_{t+1}=\underset{{\bm{H}}}{\operatorname{arg}\,\operatorname{min}}\;\quad$ $\left\lVert{\bm{H}}-{\bm{H}}_{t}\right\rVert^{2}_{\bm{W}}$ $\rhd{\bm{H}}_{t+1}\ \text{close to}\ {\bm{H}}_{t}$ (10)
@@
1.2 Validation of adherence to BFGS
@@
2 More Ablation Study and Analysis
2.1 Ablation on Incept-Mixer
@@
2.2 More visualization results
@@
2.3 Iterative results visualization
@@
2.4 More OOD results and visualization
@@
In our initial analysis, we assess the robustness of methods to out-of-distribution (OOD) data using complete patient images, computing SSIM and PSNR metrics for the entire image. While achieving the best performance across all views, we observe a degradation in performance for all methods when focusing on the circle region, as expected due to its complexity.
To address this, we extend our evaluation to specific regions, those containing the white circles. This targeted approach isolates the reconstruction performance exclusively to the circle region. Our experiments involve a randomly selected $5$ samples from the AAPM test set depicted in Fig. 13. The overall performance across the complete set of $214$ patient images is summarized in Tab. 8.
Our method significantly outperforms the second-best across all views in both SSIM and PSNR. For the most challenging case of $32$ views, we surpass the second best by $+6.47\%$ and $+1.49$ dB. With $64$ views, our performance exceeds the second best by $+10.12\%$ and $+4.21$ dB. In the easiest case of $128$ views, we outperform the second best by $+5.29\%$ and $+5.76$ dB. As anticipated, all methods exhibit degraded performance when focusing on the circle region, and the gap between our method and the second-best widens compared to the complete image. Moreover, the numerical results in Tab. 8 align with the visualizations in Fig. 13, reinforcing the robustness of our method in reconstructing abnormal data.
2.5 Reconstruction error visualization
@@
2.6 Noise Power Spectrum analysis

### v3

@@ 2.1 Background
${\bm{y}}={\bm{A}}{\bm{x}},$ (1)
where ${\bm{x}}\in\mathbb{R}^{n}$ is the (unknown) object to reconstruct with $n=h\times w$, ${\bm{y}}\in\mathbb{R}^{m}$ is the data (i.e. sinogram), where $m=n_{v}\times n_{d}$, $n_{v}$ and $n_{d}$ denote the number of projection views and detectors, respectively. ${\bm{A}}\in\mathbb{R}^{n\times m}$ is the forward model (i.e. discrete Radon transform [40]). The goal of CT image reconstruction is to recover the (unknown) object, ${\bm{x}}$, from the observed data ${\bm{y}}$. As the problem is ill-posed due to the missing data, the linear system in Eq. 1 becomes underdetermined and may have infinite solutions. Hence, reconstructed images suffer from artifacts, blurring, and noise. To address this issue, iterative reconstruction algorithms are utilized to minimize a regularized objective function with a $L^{2}$ norm constraint:
@@ 2.2 Post-processing Methods
Self-supervised learning.  SSL methods [17, 5, 58, 48, 26], have been applied for CT reconstruction. For instance, [5] proposed an equivariant imaging paradigm through a training strategy that enforces measurement consistency and equivariance conditions. To ensure equitable comparisons, we focus on supervised methods in this work.
@@ 2.3 Advancements in Deep Unrolling Networks
       ${\bm{s}}_{t}\leftarrow-{\bm{H}}_{t}\nabla_{{\bm{x}}}J({\bm{x}}_{t})$
      ${\bm{x}}_{t+1}\leftarrow{\bm{x}}_{t}+{\bm{s}}_{t}$
      ${\bm{z}}_{t}\leftarrow\nabla_{{\bm{x}}}J({\bm{x}}_{t+1})-\nabla_{{\bm{x}}}J({\bm{x}}_{t})$
      $\rho_{t}\leftarrow 1/({\bm{z}}_{t}^{\mathsf{T}}{\bm{s}}_{t})$
      ${\bm{H}}_{t+1}\leftarrow({\bm{I}}-\rho_{t}{\bm{s}}_{t}{\bm{z}}_{t}^{\mathsf{T}}){\bm{H}}_{t}({\bm{I}}-\rho_{t}{\bm{z}}_{t}{\bm{s}}_{t}^{\mathsf{T}})+\rho_{t}{\bm{s}}_{t}{\bm{s}}_{t}^{\mathsf{T}}$
@@ 3.2 Regularization term: Incept-Mixer
Figure 3: Architecture of our regularization block. It is referred to as “Incept-Mixer" and denoted as ${\mathcal{G}}$ in Eq. 4
@@
$\mathrm{Mixer}({\bm{e}}_{t})=\mathrm{Mix}(\mathrm{MLP}_{\texttt{c}},\mathrm{Mix}(\left[\mathrm{MLP}_{\texttt{h}},\mathrm{MLP}_{\texttt{w}}\right],{\bm{e}}_{t}),$ (6)
where $\mathrm{Mix}(\mathrm{Layer},{\bm{e}}_{t})=\mathrm{Layer}(\mathrm{LN}({\bm{e}}_{t}))+{\bm{e}}_{t}$, with $\mathrm{LN}$ as layer normalization. $\mathrm{MLP}_{\texttt{h}}$, $\mathrm{MLP}_{\texttt{w}}$ are applied to height and width features, respectively, and $\mathrm{MLP}_{\texttt{c}}$ to rows and shared. Finally, after $N$ such mixer layers, the regularized sample is transformed back to an image through a patch expansion step to obtain ${\mathcal{G}}({\bm{x}}_{t})$. Consequently, the iterative optimization solution is as follows:
@@ 3.4 Proposed algorithm of QN-Mixer
       ${\bm{s}}_{t}\leftarrow-{\bm{H}}_{t}{\bm{r}}_{t}$
      ${\bm{x}}_{t+1}\leftarrow{\bm{x}}_{t}+\mathcal{D}({\bm{s}}_{t})$
      ${\bm{r}}_{t+1}\leftarrow\mathcal{E}(\nabla_{{\bm{x}}}J({\bm{x}}_{t+1}))$
      ${\bm{z}}_{t}\leftarrow{\bm{r}}_{t+1}-{\bm{r}}_{t}$
      $\rho_{t}\leftarrow 1/({\bm{z}}_{t}^{\mathsf{T}}{\bm{s}}_{t})$
      ${\bm{H}}_{t+1}\leftarrow({\bm{I}}-\rho_{t}{\bm{s}}_{t}{\bm{z}}_{t}^{\mathsf{T}}){\bm{H}}_{t}({\bm{I}}-\rho_{t}{\bm{z}}_{t}{\bm{s}}_{t}^{\mathsf{T}})+\rho_{t}{\bm{s}}_{t}{\bm{s}}_{t}^{\mathsf{T}}$
@@ 4.1 Experimental Setup
[…]
Evaluation metrics.  Following established evaluation protocols [1, 49, 54], we employ the structural similarity index measure (SSIM) with parameters: level $5$, a Gaussian kernel of size $11$, and standard deviation $1.5$, as our primary performance metric. Furthermore, we supplement our assessment with the peak signal-to-noise ratio (PSNR).
@@ 4.2 Comparison with state-of-the-art methods
We evaluate frozen model performance on CT images featuring a randomly positioned white circle with noise-free sinograms, as illustrated in the third row of Fig. 4. The rationale and details are provided in the supplementary material. In Tab. 3, QN-Mixer attains state-of-the-art results across all $n_{v}$ views. First-order unrolling networks such as LEARN and RegFormer exhibit significant PSNR degradation of $-3.1$ dB and $-4.22$ dB, respectively, for $n_{v}{=}32$, while our method demonstrates a milder degradation of $-2.67$ dB.
Visual comparison.  As it can be seen on Fig. 4 and Fig. 5, FBPConvNet and DuDoTrans show significant blurring and pronounced artifacts when $n_{v}{=}32$. While Learned PD and LEARN show satisfactory performance, they struggle with intricate details, like in the liver and spine. In contrast, RegFormer produces high-quality images but faces challenges in generalizing to OOD data. QN-Mixer excels in producing high-quality images with fine details, even under challenging conditions such as $n_{v}{=}32$ views and OOD data.
Method #Iters Epoch time (s) Time (ms) #Params (M) Memory (GB)
Post-processing based denoising
FBPConvNet [19] - 68 12.4 31.1 1.30
DuDoTrans [49] - 92 60.1 15.0 1.38
First-order unrolling reconstruction networks
Learned PD [1] 10 82 47.2 0.25 0.81
LEARN [7] 30 780 679.8 4.50 1.85
RegFormer [54] 18 700 598.9 5.00 10.19
Second-order unrolling Quasi-Newton
QN-Mixer (ours) 14 594 610.2 8.50 7.83
@@
Hessian size PSNR $\uparrow$ SSIM $\uparrow$
$8^{2}\times 8^{2}$ 35.69 93.71
$16^{2}\times 16^{2}$ 38.11 95.31
$32^{2}\times 32^{2}$ 39.37 96.01
@@
Table 5: Ablation on the inverse Hessian approximation size.
Method PSNR $\uparrow$ SSIM $\uparrow$
QN with different learned regularization
Inception 31.65 85.28
U-Net 34.29 92.92
MLP-Mixer 36.89 93.87
Incept-Mixer 39.51 96.11
Pseudo-inverse ${\bm{A}}^{\dagger}$ vs Filtered Back Projection (FBP)
QN-Mixer+${\bm{A}}^{\dagger}$ 38.94 95.83
QN-Mixer+FBP 39.51 96.11
First vs second order Quasi-Newton (QN)
Incept-Mixer+first-order 37.45 94.25
Incept-Mixer+QN 39.51 96.11
Table 6: QN-Mixer ablation.
@@ 4.3 Ablation Study
Inverse Hessian approximation size.  The results in Tab. 5 emphasize the significant impact of the inverse Hessian approximation size on our performance. When too small, a notable degradation is observed (e.g., $8^{2}\times 8^{2}$), while larger sizes result in performance improvements as the approximation approaches the full inverse Hessian. Exceeding $64^{2}\times 64^{2}$ was unfeasible in our experiments due to memory constraints.
Figure 6: Ablation on the number of unrolling iterations. Left: PSNR (dB); Right: SSIM (%)
@@
Regularization term.  In Tab. 6, we evaluate the impact of the regularization term in our framework. Our Incept-Mixer is compared against various learned alternatives, including the Inception block [45] and MLP-Mixer block [46]. Additionally, employing the pseudo-inverse ${\bm{A}}^{\dagger}$ instead of the FBP results in a less pronounced degradation ($-0.57$ dB and $-0.28$%), enhancing the interpretability of QN-Mixer. Finally, we test our Incept-Mixer in the first-order framework, highlighting the significance of the second-order latent BFGS approximation with a significant improvement ($+2.06$ dB and $+1.86\%$).
@@ 5 Conclusion
In this paper, we investigate the application of deep second-order unrolling networks for tackling imaging inverse problems. To this end, we introduce QN-Mixer, a quasi-Newton inspired algorithm where a latent BFGS method approximates the inverse Hessian, and our Incept-Mixer serves as the non-local learnable regularization term. Extensive experiments confirm the successful sparse-view CT reconstruction by our model, showcasing superior performance with fewer iterations than state of-the-art methods. In summary, this research offers a fresh perspective that can be applied to any iterative reconstruction algorithm. A limitation of our work is the memory requirements associated with quasi-Newton algorithm. We introduced a memory efficient alternative by projecting the gradient to a lower dimension, successfully addressing the CT reconstruction problem. However, its applicability to other inverse problems may be limited. In future work, we aim to extend our approach to handle larger Hessian sizes, broadening its application to a range of problems.
6 Acknowledgments
This work was granted access to the HPC resources of IDRIS under the allocation 2021-[AD011012741] provided by GENCI and supported by DIM Math Innov funding.
@@ 6 Acknowledgments
Appendix A Inverse Hessian approximation
@@ A.1 Approximation via optimization
${\bm{H}}_{t+1}=\underset{{\bm{H}}}{\operatorname{arg}\,\operatorname{min}}$ $\left\lVert{\bm{H}}-{\bm{H}}_{t}\right\rVert^{2}_{\bm{W}}$ $\rhd{\bm{H}}_{t+1}\ \text{close to}\ {\bm{H}}_{t}$ (10)
@@
A.2 Validation of adherence to BFGS
@@
Appendix B More Ablation Study and visualization
B.1 Ablation on Incept-Mixer
@@
B.2 More visualization results
@@
B.3 Iterative results visualization
@@
B.4 Reconstruction error visualization
@@
Appendix C More experiments
C.1 Out-of-Distribution
@@
OOD circle performance across entire image  In the main text, we assess the robustness of methods to a simple out-of-distribution scenario, where an unseen during training, white circle is inserted, and computing SSIM and PSNR metrics for the entire image. We achieved the best performance across all views in this setting (Tab. 3). We note, however that this setting mixes the inherent performance on clean data and its ability to handle unseen patterns.
OOD circle performance in anomalous region  To further isolate the capacity of models to reconstruct unseen OOD patterns (i.e., white circle), we extend our evaluation. Instead of evaluating the whole image, we compute the reconstruction performance of a crop region containing the white circle, thus isolating the reconstruction performance exclusively to the circle region.
For that, we randomly selected 5 samples from the AAPM test set depicted. The evaluated data is depicted in Fig. 14, and the overall performance across the complete set of 214 patient images is summarized in Tab. 8 for 3 views.
Our method significantly outperforms the second-best across all views in both SSIM and PSNR. For the most challenging case of $32$ views, we surpass the second best by $+6.47\%$ and $+1.49$ dB. With $64$ views, our performance exceeds the second best by $+10.12\%$ and $+4.21$ dB. In the easiest case of $128$ views, we outperform the second best by $+5.29\%$ and $+5.76$ dB. As anticipated, all methods exhibit degraded performance when focusing on the circle region, and the gap between our method and the second-best widens compared to the complete image. Moreover, the numerical results in Tab. 8 align with the visualizations in Fig. 14.
Train AAPM $n_{v}{=}32$ AAPM $n_{v}{=}128$
Test DeepLesion $n_{v}{=}32$ AAPM $n_{v}{=}32$ AAPM $n_{v}{=}64$
Method PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
DuDoTrans 28.69 73.77 25.60 51.01 32.38 82.35
LEARN 37.21 93.46 28.20 66.10 34.40 87.21
RegFormer 37.97 94.01 29.83 72.53 37.78 92.64
QN-Mixer (ours) 38.28 94.66 30.43 80.03 37.16 91.81
Table 9: Quantitative results of the reconstruction performance of OOD cases: anatomy and geometry Bold: Best, under: second best.
Anatomy and Geometry OOD.  We explore more complex out-of-distribution (OOD) scenarios, particularly focusing on changes in anatomy and geometry. For the geometry aspect, we train our methods using AAPM data and evaluate them on DeepLesion, maintaining a fixed geometry parameter of $n_{v}{=}32$. Concerning anatomy, we test under the most challenging scenario, which involves training on AAPM data with $n_{v}{=}128$ and subsequently testing on AAPM data with two lower geometries: $n_{v}{=}32$ and $n_{v}{=}64$. The outcomes of this experiment are detailed in Tab. 9.
  • anatomy: we outperform RegFormer by $+0.31$ dB and $+0.65\%$.
  • geometry $n_{v}{=}32$: we outperform Regformer by $+0.60$ dB and $+7.50\%$ in most challenging setting.
  • geometry $n_{v}{=}64$: we lag behind Regformer by $-0.62$ dB and $-0.83\%$ in the easiest setting.
C.2 Results when A is the $90^{\circ}$ limited view CT
To further evaluate the robustness of our method, we conducted experiments on a more challenging CT inverse problem, specifically involving a $90^{\circ}$ limited-angle CT setup. In this scenario, the system matrix $A$ comprises a restricted number of views, resulting in significant artifacts and distortions in the reconstructed images. We compared the performance of our proposed QN-Mixer, with that of RegFormer on this demanding task. Our method demonstrated superior performance, achieving promising results with a PSNR of $30.17$ dB and a SSIM of $90.01\%$. This outperformed RegFormer by $+1.03$ dB in PSNR and $+0.45\%$ in SSIM. Visual results illustrating the effectiveness of our approach are provided in Fig. 15.
Figure 15: Visual comparaison on AAPM with Limited-angle CT. The system matrix $A$ is a $90^{\circ}$ limited view. The display window is set to $\left[-1000,800\right]$ HU.
C.3 MBIR evaluation
To demonstrate the superiority of our method, we conducted an additional comparison with the state-of-the-art model-based iterative reconstruction (MBIR) technique, which is widely employed in clinical settings. Our evaluation involved testing our approach against MBIR using datasets with varying numbers of views, specifically $n_{v}{=}32,64$, and $128$, with low noise added to sinograms. The corresponding PSNR values obtained with MBIR were $22.94$, $28.80$, and $34.33$ dB, with SSIM values of $67.70\%$, $72.60\%$, and $80.68\%$, respectively.
Our method, QN-Mixer, consistently outperformed MBIR across all views, showcasing an average improvement of $+12.72$ dB in PSNR and $+23.25\%$ in SSIM. These results underscore the robustness and effectiveness of our approach compared to MBIR, a benchmark technique widely utilized in clinical practice.
C.4 Noise Power Spectrum analysis